A uniform circular disc of mass ‘m’ and radius ‘a’ can rotate in a vertical plane about a horizontal axis through a point O on its circumference. The disc is held with the diameter OA through O horizontal and then released. If a constant frictional couple of 'mga/2pie' opposes the motion, the angular speed when OA is vertical will be 

Grade:12

1 Answers

Chetan Mandayam Nayakar
13 years ago

moment of inertia=ma2/2 +ma2=3ma2/2, kinetic energy=mga-((mga/2pie)*(pie/2))=(I/2)w2, where I is moment of inertia and w is angular speed.